The width of the rectangle is one-four5 of its length. If the perimeter is 40 cm, find the dimensions of the rectangle. (2 variables and full solution please, thanks!)
Answer by Boreal(15235)

width=x
length=y
half the perimeter is x+y=20
but x=(1/4)y
so y+(1/4)y=20
(5/4)y=20
y=20(4/5)=16 cm length
width is (1/4)16=4 cm width
Perimeter is 2x+2y=40 cm